OpenStudy (anonymous):

Express the quotient in simplest form 5/8 ÷15/16 please explain

OpenStudy (anonymous):

\[\frac{5}{8}\div\frac{15}{16}=\frac{5}{8}\times \frac{16}{15}\] cancel first, multiply last

OpenStudy (anonymous):

\[\frac{5}{8}\times \frac{16}{15}=\frac{\cancel{5}^1}{\cancel{8}^1}\times \frac{\cancel{16}^2}{\cancel{15}^3}=\frac{2}{3}\]

OpenStudy (anonymous):

Thank you satellite73